A non uniform rod X y of mass 40kg and length 5 m lies on the horizontal ground. It's center of gravity is from the end X. Calculate the vertical force p that will be sufficient to lift the end y from the ground why would the force p,if applied at the end X not be sufficient to lift the X from the ground
1
Expert's answer
2020-04-27T09:59:33-0400

Taking the moments about the end X:


"PL-mg(d)=0"

"P(5)-(40)(9.8)(2)=0"

"P=157\ N."

If force P was applied at the end X. Taking the moments about the end Y:


"M=PL-mg(L-d)"

"M=157(5)-(40)(9.8)(5-2)=-391<0"